The table below provides information on the price of three sampled stock, A, B, and C. Requirement: (a) using the Dow Jones Industrial Average (DJIA) price-weighted series you are to determine the index value of the stock. (b) If there is a stock split of three -for- one in stock A what will be the new divisor?

Prices of stock Before the split ($) Prices of stock after the split by stock

Prices of stock Before the split ($) Prices of stock after the split by stock A ($)
A 60 20
B 50 50
C 40 40

(c) Given a divisor of 3: (i) what will be the index value assuming there is 10% increase in the price of stock A. (ii) what will be the index value assuming there is 10% increase in the price of stock C. iii. What account for the differences in the index value due to the price changes?
Answer #1

First of All, we need to understand the actual meaning of divisor- The Divisor is the number that will give you the exact index value as it was if the stock split or dividend did not come into effect.

A) Let first calculate the Original Divisor

Name of Stock Value
A 60
B 50
C 40
SUM OF THE ABOVE 150
NO OF SHARE 3
INDEX VALUE 150/3
DIVISOR 3

B) NEW DIVISOR AFTER SHARE SPLIT

NAME OF THE SHARE PRICE
A 20
B 50
C 40
SUM OF THE ABOVE 110
OLD INDEX 50
NEW DIVISOR 110/50=2.2

C) I) INDEX VALUE IF PRICE OF STOCK OF A & C CHANGES

NAME OF THE SHARE VALUE
A 22
B 50
C 36
SUM OF THE ABOVE 108
DIVISOR 3
VALUE OF INDEX 108/3=36

iii) DIFFERENCE IN INDEX VALUE

OLD INDEX=50

NEW INDEX=36

CHANGES=50-36=14